OpenStudy (anonymous):

Help please? 3(x - 4) - 16 = 2(x - 8) + 6

OpenStudy (uri):

3(x - 4) - 16 = 2(x - 8) + 6 3x-12-16=2x-16+6 3x-2x=-16+6+12+16 x=18

OpenStudy (anonymous):

Thank you.